An ideal gas at temperature 300 K is contained under the weightless piston in a vertical cylinder. The piston area is 200 cm^2 and the atmospheric pressure is 10^5 Pa. The gas in the cylinder is heated and the piston is loaded with a weight of mass 10 kg. Find the change in gas temperature if the piston position has not changed. The answer should be expressed in kelvin [K]. Let the free fall acceleration be 10 m/s^2.
1
Expert's answer
2020-11-05T10:33:56-0500

Answer

T1=300K

P1=105pa

So P2 can be calculated as

P2=F/A+Patm


F/A="\frac{mg}{A}=\frac{100}{0.02}=5000Pa"

"P_2=0.05\times10^5+10^5=1.05\times10^5Pa"

Volume is constant so

"\frac{P}{T}=constant"


"\frac{P_1}{T_1}=\frac{P_2}{T_2}"


"\frac{10^5}{300}=\frac{1.05\times10^5}{T_2}"


"T_2=315K"